OpenStudy (anonymous):

HELPPP! Part 1: Explain, in complete sentences, the difference between a radical equation and an equation simply containing a radical. Part 2: Provide an example of each.

OpenStudy (anonymous):

Yup

OpenStudy (saifoo.khan):

from how long you havn't been doing your hw? -just wondering..

OpenStudy (anonymous):

I do all my homework.

OpenStudy (anonymous):

Oh how long I have been doing my homework? Is that what you are asking?

OpenStudy (turingtest):

lol a radical equation has a radical sign on the variable, where as if the radical is only on a constant the equation just contains it. \[\sqrt{x}+2\] IS a radical eqn.\[x+\sqrt{2}=6\] is NOT a radical eqn.

OpenStudy (turingtest):

i meant sqrt(x)+2=6 for the first eqn.

OpenStudy (anonymous):

Thank you very much, I had the answer similar to that, but I kept think it was wrong lol, I put this: The difference between a radical equation and an quation the contains a radical is that with a radical equation it can have radicals with a variable inside the radicand while an equation that contains a radical only has a number, and no variable. and for the first example: ____ √x + 2 = 7

OpenStudy (turingtest):

ah yes, quite close but, as with all mathematics, the devil is in the detail :)